# WaveRoRA: Wavelet Rotary Route Attention for Multivariate Time Series Forecasting

### v1

@@ V-D Main Results
Table II presents all results of MTS forecasting. Compared with the current SOTA Transformer-based model, namely iTransformer, WaveRoRA reduces MSE by 5.53% and MAE by 3.22%. Among all the datasets, WaveRoRA achieves the most improvement on Traffic and Electricity, which contain large amount of variables and exhibit more pronounced periodic characteristics. In particular, the reductions of MSE and MAE come to 7.82% and 3.68% on Traffic and 8.75% and 3.62% on Electricity respectively. Such degree of improvement can be attributed to two aspects: 1) The wavelet transform can accurately model varying periodic patterns over time; 2) The proposed RoRA prunes redundant inter-variable dependencies and retain key information from variables with genuine interactions.
For ETT datasets with less relevancy and non-significant periodicity, capturing inter-series dependencies may negatively impact the predicting accuracy. Models adopting the channel-independent strategy usually achieve better results, such as PatchTST and TimeNet. However, WaveRoRA shows comparable performance to PatchTST on ETTh1 and ETTm1, while achieving more accurate predictions on ETTh2 and ETTm2. This indicates that wavelet transform can effectively preserve time-frequency characteristics and approximate the benefits of time-domain modeling. In addition, we believe that high-quality relative positional information facilitates enhancing interactions between highly correlated variables while weakening those between less correlated ones.
@@ V-E Ablation Study
To justify the effectiveness of our designs, we conduct additional experiments on Weather, Traffic and Electricity. These datasets contains more series data and exhibit better stability. The results are shown in Table III.
@@ V-E1 Wavelet Learning Framework
We replace RoRA with other existing backbones to capture inter-series dependencies, which corresponds to a) w/ SA which uses Softmax attention. Since the wavelet domain embeddings are also generated through an MLP structure, w/ SA can be viewed as replacing the temporal features of iTransformer with wavelet domain features. Compared with iTransformer, w/ SA reduces the MSE by and MAE by on average, which indicates that the wavelet learning framework could provide better predicting results.
@@ V-E2 RoRA structure
We then conduct c) w/o Ro which removes the rotary positional embeddings, d) w/o Gate which removes the $\mathrm{SiLU}$-gating branch and e) w/o skip which removes the skip connection design to evaluate the components of RoRA. After these modules are removed, the predictive performance of the model decreases to varying degrees. The performance decline is most significant in w/o Gate, especially on the Traffic and Electricity datasets, where the MSE and MAE increase by an average of 4.78% and 2.83%. While the routing tokens contribute to prune redundant information, the gating unit further filter out the interaction between weakly correlated variables. The skip connection module preserve historical features and facilitates faster model convergence, therefore, w/o skip gets degraded performance.
V-F Model Efficiency
@@
As discussed in section IV-D, RoRA has linear complexity $O(MrD)$ relative to the variable number and the token dimension. To visually demonstrate the efficiency of RoRA, we set batch size to 32 on Traffic and record the training speed and GPU memory usage for WaveRoRA, iTransformer, PatchTST, Autoformer, Crossformer and DLinear. The results are shown in Fig 4, WaveRoRA achieves a good balance among predicting performance, training speed and memory usage.
@@ VI Conclusion
In this paper, we investigate to extract the underlying patterns of MTS data in the wavelet domain. We propose a wavelet learning framework to map the input MTS to wavelet embeddings and capture the inter-series dependencies with a novelly designed RoRA mechanism. The embeddings in the wavelet domain could leverage the strengths of both time and frequency domains, modeling the trends and periodic changing patterns of the time series. What’s more, the proposed RoRA introduces a small number of routing tokens to aggregate information from the $KV$ matrices and redistribute it to the $Q$ matrix. Such designs retain the powerful representational capabilities of Softmax and achieve linear complexity relative to both sequence length and token dimension. We conduct extensive experiments and WaveRoRA achieves superior performance compared to SOTA methods. Currently, our wavelet learning framework simply fits for capturing inter-series dependencies, future researches will be done to explore the combination of wavelet domain features with other types of backbones. In addition, we are going to try applying WaveRoRA on more kinds of applications such as Large Time Series Model[38] and AIOPS[37].

### v2

@@ V-D Main Results
Table II presents all results of MTS forecasting. Compared with the current SOTA Transformer-based model, namely iTransformer, WaveRoRA reduces MSE by 5.91% and MAE by 3.50%. Among all the datasets, WaveRoRA achieves the most improvement on Traffic and Electricity, which contain large amount of variables and exhibit more pronounced periodic characteristics. In particular, the reductions of MSE and MAE come to 7.94% and 4.66% on Traffic and 9.17% and 3.99% on Electricity respectively. Such degree of improvement can be attributed to two aspects: (a) The wavelet transform can accurately model varying periodic patterns over time; (b) The proposed RoRA prunes redundant inter-series dependencies and retain key information from variables with genuine interactions.
For ETT datasets with less relevancy and non-significant periodicity, capturing inter-series dependencies may negatively impact the predicting accuracy. Models adopting the channel-independent strategy usually achieve better results, such as PatchTST and TimeNet. However, WaveRoRA also shows comparable performance on ETTh1 and ETTm1, while achieving more accurate predictions on ETTh2 and ETTm2. Compared to PatchTST, the average improvements of MSE and MAE are 2.36% and 1.39%. This indicates that wavelet transform can effectively preserve time-frequency characteristics while retaining the benefits of time-domain modeling. In addition, we believe that high-quality relative positional information facilitates enhancing interactions between highly correlated variables while weakening those between less correlated ones.
@@ V-E Ablation Study
To justify the effectiveness of our designs, we conduct additional experiments on Traffic, Electricity, ETTh1 and ETTh2 and accord the average MSE and MAE. The results are shown in Table III.
@@ V-E1 Wavelet Learning Framework
We replace RoRA with other existing backbones to capture inter-series dependencies, which corresponds to a) w/ SA which uses Softmax attention. Since the wavelet domain embeddings are also generated through an MLP structure, w/ SA can be viewed as replacing the temporal features of iTransformer with wavelet domain features. Compared with iTransformer, w/ SA reduces the MSE by 4.40% and MAE by 1.36% on average, which indicates that the wavelet learning framework could provide better predicting results.
Figure 5: The MSE results with prediction length $H=96$ on ETTh1 and ETTh2. Varying settings of DWT decomposition level $J$ and dropout values are applied. The model’s prediction accuracy improves and stabilizes with a larger $J$, while different dropout values have a minimal effect on performance.
@@ V-E2 RoRA structure
We then conduct c) w/o Ro which removes the rotary positional embeddings and d) w/o skip which removes the skip connection design to evaluate the components of RoRA. After these modules are removed, the predictive performance of the model decreases to varying degrees. The performance decline is most significant in w/o Gate on the Traffic and Electricity datasets, where the MSE and MAE increase by 5.03% and 3.81% on average. However, for ETTh1 and ETTh2, removing the skip connection layer causes more deterioration in the prediction accuracy, where the MSE and MAE increase by an average of 4.25% and 1.95%. For datasets with a larger number of variables, the feature diversity carried by series-wise tokens is generally guaranteed, making it beneficial to select high-quality information for more accurate predictions. In contrast, for datasets with fewer variables, enhancing the feature diversity often leads to improved performance.
V-F Hyper-parameter Sensitivity Analysis
V-F1 DWT decomposition levels & Dropout
DWT hierarchically decomposes a time series into low-pass and high-pass components, corresponding to coarse trends and finer periodical details at different scales. With the increase of decomposition level $J$, coarser periodic information is separated and the low-pass component will exhibit more significant trend characteristics. To obtain more comprehensive results, we change the value of dropout layers at the same time. Dropout layers improve the generalization ability of the model by randomly masking part of the input. A proper dropout value is conducive to avoiding overfitting. The dropout layer mainly works in each WaveNorm layer. We set $J=\{0,1,2,3,4,5\}$ and $dropout=\{0.0,0.1,0.2,0.3\}$. Note that when $J=0$, the model does not make any DWT decomposition and RoRA is applied to time-domain features. We conduct the experiments on ETTh1 and ETTh2 and the final results are shown in Fig 5. Generally, different dropout values do not cause substantial changes of model performance. In contrast, whether to operate DWT to use wavelet-domain features could lead to significant variations in the results. Specifically, when $J=0$, i.e. only time-domain features are used, the model provides suboptimal results. With the increase of $J$, the model’s prediction accuracy improves and stabilizes. This indicates that key periodic components is separated and the remaining trend components no longer contain extra periodic information.
V-F2 Number of routing tokens
Typically, the routing tokens are kept with $r\ll M$. A larger $r$ could lead to more comprehensive capturing of element-wise correlation but may also cause more information redundancy. To evaluate the effect of $r$, we manually set $r=\{2,20,40,200\}$ and record the average MSE and MAE on Weather, Electricity and ETTh2. The settings roughly cover the range of variable numbers of these datasets. Although a larger number of router tokens may intuitively aggregate information from the $K$$V$ matrices more comprehensively, the results in Fig LABEL:fig:r provide limited support for this assumption. For Electricity, the best $r$ is 10 or 20, while $r=2$ yields best performance on ETTh2. This may be due to the potentially sparse dependencies between variables. Specifically, while complex associations may exist among multivariate data, the patterns of the correlations are likely to be similar. Therefore, a small number of routing tokens could also filter out the redundant information, enhancing the model to focus on more critical inter-series dependencies. Nonetheless, the model performance the overall model performance does not exhibit substantial fluctuations with varying $r$, which demonstrates the robustness of our proposed model.
V-G Effect of Varying Wavelet Basis
Unlike the Fourier transform, the results of wavelet transform vary greatly depending on the chosen wavelet function $\psi$ and scale function $\phi$. Therefore, to explore how different wavelet bases impact model prediction performance across various types of datasets, we selected three wavelet bases: Haar wavelet, Symlet wavelet and Coiflet wavelet[22]. The Haar wavelet, one of the earliest orthogonal wavelet functions with compact support, is the simplest wavelet function used in wavelet analysis. While computationally efficient, it has limited performance as a wavelet basis due to its discontinuity in the time domain. The Symlet wavelet is a nearly symmetric wavelet function that improves the Daubechies function. It reduces phase distortion to some extent during signal analysis and reconstruction. Generally, the Symlet wavelet is more beneficial for signals with stronger periodic characteristics. In this paper, we primarily use the Sym3 wavelet for DWT. The Coiflet wavelet, which is known for its higher symmetry, could simultaneously maintain high vanishing moments on its wavelet function and scale function. It usually performs well in handling signals with abrupt features. In this experiment, we use Coif3 to evaluate the effectiveness of Coiflet wavelet. The results can be seen in Fig LABEL:fig:basis. For Traffic and Electricity that exhibit more pronounced periodic characteristics, the Sym3 wavelet leads to the best predicting performance, while for ETTh1, the best wavelet function changes to Coif3. This indicates that different wavelet functions can provide tokens with varying quality when handling data with different characteristics.
V-H Model Efficiency
@@
As discussed in section IV-D1, RoRA has linear complexity $O(MrD)$ relative to the variable number and the token dimension. To visually evaluate the efficiency of WaveRoRA and other baseline models, we set $r=20$ and batch size $B=32$ on Traffic and record the training speed and GPU memory usage for WaveRoRA, iTransformer, PatchTST, Autoformer, Crossformer and DLinear. The results are shown in Fig 7. WaveRoRA achieves a good balance among predicting performance, training speed and memory usage.
@@ VI Conclusion
In this paper, we investigate to extract the underlying patterns of MTS data in the wavelet domain. We propose a wavelet learning framework to map the input MTS to wavelet embeddings and capture the inter-series dependencies with a novelly designed RoRA mechanism. The embeddings in the wavelet domain could leverage the strengths of both time and frequency domains, modeling the trends and periodic changing patterns of the time series. What’s more, the proposed RoRA introduces a small number of routing tokens to aggregate information from the $KV$ matrices and redistribute it to the $Q$ matrix. Such designs retain the powerful expressivity of Softmax and achieve linear complexity relative to both sequence length and token dimension. We conduct extensive experiments and WaveRoRA achieves superior performance compared to SOTA methods. Currently, our wavelet learning framework simply fits for models capturing inter-series dependencies. What’s more, different types of wavelet functions exhibit potentials for processing MTS data with varying patterns. Future researches will be done to (a) explore the combination of wavelet domain features with other types of backbones and (b) find an automatic approach to adaptively select the best wavelet function for specific data. In addition, we are going to try applying WaveRoRA on more kinds of applications such as Large Time Series Model[40] and AIOPS[38].